Find a polynomial with integer coefficients that satisfies the given conditions.
Q(X) has degree 3 and zeros −7 and 1 + i.

Answer by edjones(8007)

x=-7, x=1+i, x=1-i
(x+7)(x-1-i)(x-1+i)
=x^3+5x^2-12x+14
.
Ed